# Beyond Markovian: Reflective Exploration via Bayes-Adaptive RL for LLM Reasoning

### v1

@@ B.1 Evaluation Results: Accuracies
In this section, we provide the evaluation accuracy results for Qwen2.5-Math-1.5B, Qwen2.5-Math-7B, and DeepSeek-R1-Distill-Llama-8B in Figure 9, 10, and 11, respectively. In addition to the benchmark scores in Table 1, we also report the performance of the models on AIME 2024 and AMC 2023. It can be observed that BARL outperforms Markovian RL baselines in terms of both accuracy and convergence rate on most of the reported benchmarks. Again, the performance on AIME and AMC benchmarks may be further enhanced by choosing harder training data with increased response length, especially for R1-Distill-Llama-8B fine-tuned models whose initial average response lengths on AIME and AMC ($2008$ and $1886$, respectively) exceed the maximum training length ($1024$).
@@ B.3 Token Efficiency without Greedy Decoding Outputs
Figure 16: Ablation on token efficiency and pass@k accuracies with sampling temperature$=1$. GRPO and the base models are less robust to temperatures.
@@ B.4 Some Unsuccessful Attempts
As a straightforward implementation of the Bayes-Adaptive RL policy gradient in (5.1), we explored using value ensembles to estimate the posterior-weighted value $\mathbb{E}_{\mathcal{M}\sim p(\mathcal{M}|h_{t})}[Q^{\pi_{\theta}}_{\mathcal{M}}(s_{t},a_{t})]$. Specifically, we trained an ensemble of state-action value functions to capture epistemic uncertainty. We experimented with two approaches to constructing the ensemble: (1) fine-tuning multiple linear value heads on disjoint data subsets, each paired with chain-of-thought (CoT) trajectories and outcome rewards; and (2) applying Bayesian LoRA in a way similar to [48]. However, both methods failed to effectively capture epistemic uncertainty, likely because different ensembles still share a large fraction of parameters. While maintaining independent value models may better capture this uncertainty, doing so incurs substantial computational cost. We leave the development of more efficient implementations to future work.

### v2

@@ B.1 Evaluation Results: Accuracies
In this section, we provide the evaluation accuracy results for Qwen2.5-Math-1.5B, Qwen2.5-Math-7B, and DeepSeek-R1-Distill-Llama-8B in Figure 9, 10, and 11, respectively. All models are trained using the same random seed and hyperparameters as described in Section 7.1. It can be observed that BARL outperforms the two RL baselines in terms of both accuracy and convergence rate on most of the reported benchmarks.
@@ B.3 Token Efficiency without Greedy Decoding Outputs
Figure 16: Ablation on token efficiency and pass@k accuracies with sampling temperature=1=1. GRPO and the base models are less robust to temperatures.
@@ B.4 Some Unsuccessful Attempts
As a straightforward implementation of the Bayes-Adaptive RL policy gradient in (5.1), we explored using value ensembles to estimate the posterior-weighted value $\mathbb{E}_{\mathcal{M}\sim p(\mathcal{M}|h_{t})}[Q^{\pi_{\theta}}_{\mathcal{M}}(s_{t},a_{t})]$. Specifically, we trained an ensemble of state-action value functions to capture epistemic uncertainty. We experimented with two approaches to constructing the ensemble: (1) fine-tuning multiple linear value heads on disjoint data subsets, each paired with chain-of-thought (CoT) trajectories and outcome rewards; and (2) applying Bayesian LoRA. However, both methods failed to effectively capture epistemic uncertainty, likely because they fine-tune only a small subset of LLM parameters, which is insufficient to fully represent the uncertainty. While maintaining independent value models may better capture this uncertainty, doing so incurs substantial computational cost. We leave the development of more efficient implementations to future work.